OpenStudy (ffasinger):

What is the equation of the following line? Be sure to scroll down first to see all answer options.

OpenStudy (ffasinger):

A.y=1/8x b y=2x c y=4x d y=-1/4x e y=1/4x f y=1/8 x

OpenStudy (phi):

do you know how to find the slope, which is "change in y divided by change in x" between 2 points ?

OpenStudy (ffasinger):

its been a while

OpenStudy (phi):

the two points are (8,2) and (0,0) the 2nd number in each pair are the y's first do 2-0 then do 8-0 then divide the change in y (first answer) by the change in x (2nd answer)

OpenStudy (ffasinger):

ok

OpenStudy (phi):

can you find the slope?

OpenStudy (ffasinger):

i can try

OpenStudy (phi):

first do 2-0 then do 8-0 then divide the change in y (first answer) by the change in x (2nd answer)

OpenStudy (ffasinger):

i am confused

OpenStudy (ffasinger):

@Dani_Rose

OpenStudy (ltrout):

Slope is y2-y1/x2 - x1 You want to do 2-0/8-0, which comes out to 2/8 or 1/4. The equation of a line follows the pattern y = mx + b. M being the slope, b being the y-intercept. So far your equation is y = 1/4x + b. If you graph this right now, you can see that it goes through the point (8,2) and (0, 0) so you don't have a b. So your equation is y = 1/4x. Does that make sense?

OpenStudy (ffasinger):

@okdutchman7

OpenStudy (okdutchman7):

In order to find the equation you must understand the rise and the run. Do you know what they are?

OpenStudy (anthonyym):

The slope of a line is the change in y over change in x, or (y2 - y1) / (x2 - x1). Just pick two points, any two points, (x1, y1) and (x2, y2), and the slope will be (y2 - y1) / (x2 - x1).

OpenStudy (anthonyym):

The equation of a line in slope intercept form is y = mx + b, where m is slope and b is the y-intercept. See above for slope. The y-intercept is where the graph intersects the y-axis. In this case the y-intercept is (0,0). Substitute the slope and y-intercept into the equation and you will have the equation of the line.